# Automated Identification and Segmentation of Hi Sources in CRAFTS Using Deep Learning Method

### v1

@@ 5 Summary
In this work, we propose a novel method for Hi source detection that harnesses the power of 3D-Unet segmentation network to accurately identify and segment Hi sources. Experimental results demonstrate remarkable performance on our custom test set, achieving high recall (91.6%) and accuracy (95.7%), while maintaining good consistency across different datasets.
Compared to the SoFiA software, our proposed approach exhibits a significant improvement in recognition precision and attains satisfactory segmentation outcomes within the context of our dataset. Comparative analysis with state-of-the-art network architectures such as Swin-UNETR and UX-Net indicates that customizing the network architecture in accordance with the specific attributes of the data and target features is indeed a critical factor in optimizing the overall functionality and performance of the model. This not only validates the efficacy of our adopted method but also highlights the profound value of our tailored dataset in enhancing the precision and efficiency of Hi source detection tasks.
Additionally, the meticulously constructed and annotated custom dataset we have developed plays a pivotal role in future identification tasks concerning Hi sources. This dataset encompasses a rich array of Hi source examples, covering a wide range of observing conditions and signal strengths, with a particular emphasis on cases where Hi sources are difficult to discern amidst complex background noise and low signal-to-noise ratio environments. Through diligent manual annotation, we ensure the authenticity and integrity of every source in the dataset, which is essential for training and validating identification algorithms.
@@
Furthermore, given the success with our custom dataset and architecture, future directions include expanding the dataset diversity, developing adaptive learning strategies, and exploring ways to integrate extra contextual information to boost the accuracy of Hi source identification and segmentation.
@@ Acknowledgements
This work was Supported by National Key R&\&&D Program of China No. 2022YFB4501405 and National Natural Science Foundation of China grant No. 12373026.
@@
The labeled CRAFTS Data used in this paper will be available in the near future, and the data access URLs will be synchronized on GitHub https://github.com/fishszh/HISF.

### v2

@@ 5 Summary
In this work, we propose a novel method for Hi source detection that harnesses the power of 3D-Unet segmentation network to accurately identify and segment Hi sources. Experimental results demonstrate remarkable performance on our test set, achieving high recall (91.6%) and accuracy (95.7%), while maintaining good consistency across different datasets.
Compared to the SoFiA software, our proposed method demonstrates a significant improvement in recognition precision and attains satisfactory segmentation outcomes within the context of HISF dataset. It exhibits superior generalization capabilities, effectively mitigating the impact of RFI and other data processing artifacts to a certain extent. Comparative analysis with state-of-the-art network architectures such as Swin-UNETR and UX-Net indicates that customizing the network architecture in accordance with the specific attributes of the data and target features is indeed a critical factor in optimizing the overall functionality and performance of the model. This not only validates the efficacy of our adopted method but also highlights the profound value of our tailored HISF dataset in enhancing the precision and efficiency of Hi source detection tasks.
Additionally, the meticulously constructed and annotated custom HISF dataset we have developed plays a pivotal role in future identification tasks concerning Hi sources. HISF dataset offers a comprehensive collection of Hi source instances, covering a wide range of observing conditions and signal strengths, with a particular emphasis on cases where Hi sources are difficult to discern amidst complex background noise and low SNR environments. Our meticulous manual annotation process guarantees the authenticity and completeness of each source within the HISF dataset, a critical factor for the training and validation of Hi source identification algorithms.
@@
Furthermore, given the success with our custom HISF dataset and architecture, future directions include expanding the dataset diversity, developing adaptive learning strategies, and exploring ways to integrate extra contextual information to boost the accuracy of Hi source identification and segmentation.
@@ Acknowledgements
This work was Supported by National Key R&\&&D Program of China No. 2023YFE0110500 and No. 2022YFB4501405, National Natural Science Foundation of China grant No. 12373026 and the Leading Innovation and Entrepreneurship Team of Zhejiang Province of China Grant No. 2023R01008.
@@
The labeled CRAFTS Data used in this paper will be available in the near future, and the data access URLs will be synchronized on HIverse and GitHub https://github.com/fishszh/HISF.